OpenStudy (anonymous):

I need to know how to graph y=2x^2?? Can you please explain how I find the points??

OpenStudy (anonymous):

When it comes to graphing all you need to do is plug in your values and find out what it gives. There's no good feature on this site that will show you the coordinate system. But basically: If x = 0 you get a value of y = 2*0^2 = 0, So your graph starts at 0. If x = 1 you get a value of y = 2*1^2 = 2 You now have two points in your graph, which is (0,0) and (1,2), keep filling out these points with more values of x and y, and sketch your graph on a piece of paper and you'll see how the graph itself will look like. Remember to use negative values as well, if the assignment asks for it.

OpenStudy (anonymous):

Ok.. so, am I on the right track by saying that the next point would be (2, 8)? And if I plug in negative numbers would it be (-1,2) and (-2,8)??? Cause those are the same y points???

OpenStudy (anonymous):

Yes, that's correct. That's the nice part of functions that look like x^2, they all give the same value regardless if the x is negative or positive, good you saw that. :)